# Data Mesh: a Systematic Gray Literature Review

### v1

@@ 6.4.2 Policy Agent 4 S2,S29,S78,S97
The data products in a data mesh should be interoperable to enable the seamless composition of products and foster cross-domain collaboration. The data products also need to be portable so that self-serve platform services and governance applications can be modified or replaced with little or no modification to data products. The standardization of the data mesh can enable its interoperability and portability. The standards are necessary for i) defining data products, their interfaces, and their composition; ii) data product publishing and discovery, quality assessment, deployment, and governance; iii) self-serve platform services and their interfaces; and iv) artifacts used/produced by data products (e.g., data models, metadata, and ML models). The standardization efforts in the SOA and cloud domains can potentially help address these challenges.
@@
The data mesh literature provides general principles for developing a data mesh architecture, e.g., domain-driven design, product thinking, platform thinking, and computational governance. However, no tools and detailed (empirically validated) methodologies exist for applying these principles to develop and operate a data mesh. Organizations would also need guidance and tools for systematically migrating their legacy data architectures into data mesh while assessing costs and benefits.
@@
Data products are at the heart of a data mesh. Each phase of a data product life cycle (e.g., identification, design, development, testing, deployment, discovery, composition, operation, monitoring, and evolution) can have open challenges. For example, the product identification phase decomposes the analytical data landscape of an organization into a set of data products. The partitioning process needs to support applying domain-driven design correctly, ensure the modularity and utility of data products, and consider reuse of the artifacts of the existing data architecture, such as data source systems, data pipelines, machine learning pipelines, documentation, and execution logs. Further exploratory research studies are needed to properly understand the research challenges in each phase of the product life cycle. Due to the comparability of a data product and a service, we believe the existing relevant research studies on SOA and microservices can help understand and address the above product challenges.
@@
The domain teams need to be able to use domain-agnostic platform services in a self-service manner to easily and independently build and manage their data products. While there exist research studies on self-serve systems, e.g., self-service BI (business intelligence) (Lennerholt et al., 2018) and self-service portals for cloud applications (Mietzner and Leymann, 2010), they do not cover each platform service type, e.g., data pipelines, ML pipelines, and IaC based data product provisioning. Moreover, the recommender systems can also be incorporated into platform services to assist users with building data products, e.g., providing recommendations interactively when creating a data pipeline or IaC blueprint. Finally, since the organizations may already use data platforms, they would potentially need the guidance and techniques to add self-service capabilities to their legacy data platforms.
@@
While data governance has been an active research topic for several decades, as noted in several recent reviews (Abraham et al., 2019; Davidson et al., 2023; Janssen et al., 2020b), there are still many open research issues. Data mesh, especially its principles of decentralized ownership of data and federated computational governance, may add new dimensions to data governance research issues. For example, the efficient composition of data products across domain boundaries (and potentially across organizational boundaries) depends on the ability of the federated governance mechanisms to ensure that data products are interoperable, compliance with QoS (Quality of Service) and policy requirements, and in line with the company-wide data strategy. Moreover, the end-to-end data governance automation using the policy-as-code approach (i.e., computational governance) is largely unexplored.
@@
As discussed in Section 5.1, data mesh introduces a set of new roles and responsibilities for teams and individuals in an organization. In particular, the responsibilities of domain data are shifted closer to decentralized autonomous domain teams. Hence, for organizations that use centralized operating models and organizational designs, embarking data mesh journey may require a significant organizational change (in addition to technological changes). Data mesh can also significantly impact the existing data management and governance processes and practices in an organization. Thus, more studies are necessary to understand the organizational impacts of data mesh, the barriers to the adoption of data mesh, and to create guidelines and frameworks to perform and manage required organizational changes.
@@ 7.1. Studies on Data Mesh
Several studies on the application of data mesh architectures exist. Loukiala et. al. describe migrating from centralized, monolithic data platform architectures to a decentralized data-mesh-like architecture at a large Nordic manufacturing country (Loukiala et al., 2021). Their work emphasizes concrete steps that can be taken during such a migration and provides high-level contrasts between traditional data warehouse architecture and data mesh. Machado et. al. present several works on data mesh anatomy and architecture. The first is an overview and explanation of data mesh concepts highlighted by two implementations in the industry (an online retailer, i.e., Zalando, and a streaming platform, i.e., Netflix) (Araújo Machado et al., 2022). Then, in another paper, they introduce a “domain model” that contains the abstract classes of a data mesh and a high-level architecture (Machado et al., 2021). Finally, based on these works, they investigate existing (commercial) tools that can be used to realize a data mesh implementation and validate their architectural model through a proof-of-concept implementation demo (Machado et al., 2022). Similarly, Butte and Butte (Butte and Butte, 2022) present a more-complete reference architecture. However, it is unclear what their architecture is based on, as the paper provides no mThe clearinghouse logs data transactions we know, these are currently the only published academic works explicitly discussing data mesh architecture. However, their descriptions of the data mesh concepts lack a thorough exploration of the state-of-the-art in data mesh. Instead, they are almost exclusively based on the blog posts of Zhamak Dehghani (Dehghani, 2020; (2019), Thoughtworks), as well as two corporate presentations: Zalando (Max and Arif, 2020) and Netflix (Cunningham, 2020).
@@ 8. Threats to Validity
This section discusses the potential threats to external, construct, and internal validity (Wohlin et al., 2012) that may apply to our study.
First, we note the risk of missing relevant studies for the systematic gray literature review. Since the field is developing rapidly, it can be interesting to repeat the gray literature review after some time to see how the topic matures as it is adopted more by the industry. For this reason, we extended the study to include sources from September 2021 to May 2022, which yielded 35 new sources.
Additionally, as part of the selection process of our methodology, we rely on the (subjective) expertise of the authors to determine the quality and relevance of a source. Such a process runs the risk of introducing bias to the final selection. Therefore, two authors assessed each source, and an inter-reliability kappa coefficient was calculated to ensure bias was not too high.
Finally, the reference architecture was created systematically by looking at various sources. However, as was noted above, few complete, holistic architecture instances exist to compare with, and many sources only consider partial architectures. These partial architecture instances might lead to a sub-optimal holistic reference architecture. At the same time, we note that this problem is somewhat cyclical: a reference architecture can contribute significantly to the number of complete architecture instantiations, which in turn can improve future reference architectures.
@@ 9. Conclusion and Future Work
Although enterprises are increasingly becoming data-driven, their centralized monolithic data architectures still prevent them from leveraging the full value of their analytical data. The data mesh is an emerging domain-driven decentralized architecture and sociotechnical paradigm that promises to address the limitations of centralized enterprise data architectures. Even though the data mesh is gaining significant attention in the industry, there is an evident scarcity of academic research on the topic. This paper systematically selects and reviews 114 industrial gray literature on data mesh to identify key design principles, architectural components, organizational roles, benefits, and concerns. The findings from the review were further enriched with the academic literature from similar domains to create reference architectures for its key dimensions. Finally, we identified several research challenges that need to be addressed to realize the vision of the data mesh. We believe the identified open issues and reference architectures can provide a framework for future research in the data mesh domain.
We are currently working on domain-specific languages for defining data products and architecture decision-marking frameworks for three pillars of the data mesh: data as a product, self-serve platforms, and federated governance. Furthermore, we plan to develop a methodology for designing and instantiating a data mesh architecture by leveraging the reference architectures in the paper.

### v2

@@ 5.3. A Logical Runtime Structure of Data Mesh
The data products in a data mesh should be interoperable to enable the seamless composition of products and foster cross-domain collaboration. The data products also need to be portable so that self-serve platform services and governance applications can be modified or replaced with little or no modification to data products. The standardization of the data mesh can enable its interoperability and portability (AgileLab, 2021; Rajagopalan, 2021; Dehghani, 2019a; Bauer, 2021; Jochen Christ and Harrer, 2022; Strengholt, 2022c). The standards are necessary for i) defining data products, their interfaces, and their composition; ii) data product publishing and discovery, quality assessment, deployment, and governance; iii) self-serve platform services and their interfaces; and iv) artifacts used/produced by data products (e.g., data models, metadata, and ML models). The standardization efforts in the SOA and cloud domains can potentially help address these challenges. For example, the Web Service Description Language (WSDL) supports modeling services with diverse endpoints, message types, and access policies (e.g., HTTP endpoints, publish-subscribe topics, and FTP endpoints) (Papazoglou et al., 2007). Hence, WDSL and other related web service standards may help identify the concepts for modeling different types of data products. Topology and Orchestration Specification for Cloud Applications (TOSCA) is an OASIS standard language used to model the deployment topologies of complex applications (Bellendorf and Mann, 2020). It can be a suitable candidate for a technology-agnostic IaC language that can define the interoperable deployment blackprints for data products and self-serve platform components.
@@
The data mesh literature (Dehghani, 2022b; Moses, 2020; Perlov, 2021; Oracle, 2021a; Rajagopalan, 2021; Oostra, 2021b; Ramos, 2021; Strengholt, 2022c) provides general principles for developing a data mesh architecture, e.g., domain-driven design (DDD), product thinking, platform thinking, and computational governance. However, no tools and detailed (empirically validated) methodologies exist for applying these principles to develop and operate a data mesh. Organizations would also need guidance and tools for systematically migrating their legacy data architectures into data mesh while assessing costs and benefits. The research issues may benefit from the rich knowledge about applying DDD for designing and implementing microservices (Zhong et al., 2024) and migrating legacy monolithic architectures into cloud-based microservices architectures (Hasan et al., 2023).
@@
Data products are at the heart of a data mesh. Each phase of a data product life cycle (e.g., identification, design, development, testing, deployment, discovery, composition, operation, monitoring, and evolution) can have open challenges (Dehghani, 2022b; Rajagopalan, 2021; Orwat, 2021; Reid, 2021; Strenholt and Courtright, 2022; Strengholt, 2020; Dowsett, 2021; Ramos, 2021; Strengholt, 2022b). For example, the product identification phase decomposes the analytical data landscape of an organization into a set of data products. The partitioning process needs to support applying domain-driven design correctly, ensure the modularity and utility of data products, and consider reuse of the artifacts of the existing data architecture, such as data source systems, data pipelines, machine learning pipelines, documentation, and execution logs. Further exploratory research studies are needed to properly understand the research challenges in each phase of the product life cycle. Due to the comparability of a data product and a service, we believe the existing relevant research studies on SOA and microservices can help understand and address the above product challenges. For example, there exist literature on web service discovery (Ngan and Kanagasabai, 2013) and composition (Sheng et al., 2014), DDD for microservices development (Zhong et al., 2024), and metadata management techniques for data stores (Singh and Bawa, 2018) and web services (Tosi and Morasca, 2015).
@@
The domain teams need to be able to use domain-agnostic platform services in a self-service manner to easily and independently build and manage their data products. While there exist research studies on self-serve systems, e.g., self-service BI (business intelligence) (Lennerholt et al., 2018) and self-service portals for cloud applications (Mietzner and Leymann, 2010), they do not cover each platform service type, e.g., data pipelines, ML pipelines, and IaC based data product provisioning. Moreover, the recommender systems can also be incorporated into platform services to assist users with building data products, e.g., providing recommendations interactively when creating a data pipeline or IaC blackprint. Finally, since organizations may already use data platforms, they may need guidance and techniques to add self-service capabilities to their legacy data platforms (Dehghani, 2019b; Gillin, 2021; Keboola, 2021; Snowflake, 2021; Cooke, 2021; Kumar, 2022; Wolf, 2022). The existing literature on recommender systems for software engineering investigates simplifying and assisting users in each phase of the software development lifecycle (Gasparic and Janes, 2016; Almonte et al., 2022). Many studies have explored generating code, including SQL and data processing code, from natural language descriptions (Xu et al., 2022). LLMs (Large Langauge Models) have also shown great potential in this research area (Rajbhoj et al., 2024; Zhang et al., 2024).
@@
While data governance has been an active research topic for several decades, as noted in several recent reviews (Abraham et al., 2019; Davidson et al., 2023; Janssen et al., 2020), there are still many open research issues. Data mesh, especially its principles of decentralized ownership of data and federated computational governance, may add new dimensions to data governance research issues. For example, the efficient composition of data products across domain boundaries (and potentially across organizational boundaries) depends on the ability of the federated governance mechanisms to ensure that data products are interoperable, comply with QoS (Quality of Service) and policy requirements, and in line with the company-wide data strategy (Dehghani, 2022b; Perlov, 2021; Rajagopalan, 2021; Kumar, 2022; Strengholt, 2022c). Moreover, the end-to-end data governance automation using the policy-as-code approach (i.e., computational governance) is largely unexplored. The policy-as-code approach has been applied in use cases such as enforcing policies in microservice architectures (using the service mesh) (Chandramouli et al., 2020) and IaC-based deployment pipelines (Zeini et al., 2023), as well as enforcing contracts in machine learning environments (Truong and Nguyen, 2021).
@@
As discussed in Section 5.1, data mesh introduces a set of new roles and responsibilities for teams and individuals in an organization. In particular, the responsibilities of domain data are shifted closer to decentralized autonomous domain teams. Hence, for organizations that use centralized operating models and organizational designs, embarking on a data mesh journey may require a significant organizational change (in addition to technological changes). Data mesh can also significantly impact an organization’s existing data management and governance processes and practices. Thus, more studies are necessary to understand the organizational impacts of data mesh and the barriers to the adoption of data mesh, as well as to create guidelines and frameworks to perform and manage required organizational changes (Dehghani, 2019b; Gillin, 2021; Perlov, 2021; Keboola, 2021; Rajagopalan, 2021; Bauer, 2021; Strenholt and Courtright, 2022). The research on migrating on-premise systems to the cloud and legacy applications to microservices has also studied organizational challenges and potential solutions for them (Gholami et al., 2016; Vitharana and Daya, 2024). Further research could also help to understand and address the organizational challenges in data mesh transition.
@@ 7.1. Studies on Data Mesh
Loukiala et. al. describe migrating from centralized, monolithic data platform architectures to a decentralized data-mesh-like architecture at a large Nordic manufacturing country (Loukiala et al., 2021). Their work emphasizes concrete steps that can be taken during such a migration and provides high-level contrasts between traditional data warehouse architecture and data mesh. Machado et. al. present several works on data mesh anatomy and architecture. The first is an overview and explanation of data mesh concepts highlighted by two implementations in the industry (an online retailer, i.e., Zalando, and a streaming platform, i.e., Netflix) (Araújo Machado et al., 2022). Then, in another paper, they introduce a “domain model” that contains the abstract classes of a data mesh and a high-level architecture (Machado et al., 2021). Finally, based on these works, they investigate existing (commercial) tools that can be used to realize a data mesh implementation and validate their architectural model through a proof-of-concept implementation demo (Machado et al., 2022). Similarly, Butte and Butte (Butte and Butte, 2022a) present a more-complete reference architecture. However, it is unclear what their architecture is based on, as the paper provides no clear references to the sources. These are currently the only published academic works explicitly discussing data mesh architecture. However, their descriptions of the data mesh concepts lack a thorough exploration of the state-of-the-art in data mesh. Instead, they are almost exclusively based on the blog posts of Zhamak Dehghani (Dehghani, 2022b, 2019b), as well as two corporate presentations: Zalando (Max and Arif, 2020) and Netflix (Cunningham, 2020).
Kumar and Sujata (Butte and Butte, 2022b) describe the key concepts of the data mesh, present a template for a data processing pipeline within a domain, and provide an example of a data mesh deployment using AWS cloud. Li et. al. (Li et al., 2022) presents a case study of applying data mesh concepts to power system data management. They use a federated data processing engine and a metadata layer over raw data sources to build value-added data services. Podlesny et. al. (Podlesny et al., 2022) discusses the potential data privacy issues that a distributed data mesh can have, focusing on exposing PII (Personal Identifiable Information) data through quasi-identifiers. Kraska et. al. (Kraska et al., 2023) propose a visionary system that aims to provide a unified data processing interface over a federated set of data processing engines, allowing non-expert end-users to use the data to access different specialized engines. Dahdal et. al. (Dahdal et al., 2023) presents an architecture for a middleware system on the edges that offers data-intensive applications for tactical battlefields as data products. The middleware offers common data management features such as data ingestion, storage, security, and APIs. Hooshmand et. al. (Hooshmand et al., 2022) proposes applying the data mesh approach to engineering product lifecycle management (PLM). They also suggest applying domain-driven design and using ontologies for data modeling.
@@ 8. Threats to Validity
This section discusses the potential threats to external, construct, and internal validity (Wohlin et al., 2012) that may apply to our study. First, we note the risk of missing relevant studies for the systematic gray literature review. Since the field is developing rapidly, it can be interesting to repeat the gray literature review after some time to see how the topic matures as the industry adopts it more. Hence, we extended the study to include sources from September 2021 to May 2022, which yielded 35 new sources.
Additionally, as part of our methodology’s article selection process, we rely on the authors’ (subjective) expertise to determine a source’s quality and relevance. Such a process risks introducing bias to the final selection. Therefore, two authors assessed each source, and an inter-reliability kappa coefficient was calculated to ensure bias was not too high.
Finally, the reference architecture was created systematically by examining various sources. However, as noted above, few complete, holistic architecture instances exist to compare with, and many sources only consider partial architectures. These partial architecture instances might lead to a sub-optimal holistic reference architecture. At the same time, we note that this problem is somewhat cyclical: a reference architecture can contribute significantly to the number of complete architecture instantiations, which in turn can improve future reference architectures.
@@ 9. Conclusion and Future Work
Although enterprises are increasingly becoming data-driven, their centralized monolithic data architectures still prevent them from leveraging the full value of their analytical data. The data mesh is an emerging domain-driven decentralized architecture and sociotechnical paradigm that promises to address the limitations of centralized enterprise data architectures. Even though the data mesh is gaining significant attention in the industry, there is an evident scarcity of academic research on the topic. This paper systematically selects and reviews 114 industrial gray literature on data mesh to identify key design principles, architectural components, organizational roles, benefits, and concerns. The findings from the review were further enriched with the academic literature from similar domains to create reference architectures for its key dimensions. Finally, we identified research challenges that need to be addressed to realize the vision of the data mesh. We believe the identified open issues and reference architectures can provide a framework for future data mesh research.
As for future work, we will propose domain-specific languages to define data products and decision-marking platforms for the three pillars of the data mesh: data as a product, self-serve platforms, and federated governance. We will create the metamodels for data products and contracts and develop a method to design and instantiate a data mesh architecture by leveraging the reference architectures provided in this paper.